OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its for sides, If the width of a rectangle is 8inches and its perimerter is 48 in,find the length of the rectangle?

OpenStudy (anonymous):

2w+2l=p 2*8 + 2l = 48 16 + 2l = 48 2l = 48 - 16 you finish

OpenStudy (anonymous):

P = 2L + 2W 48 = 2L + 2(8) So, now, just solve for "L". Can you simplify this last equation?